# Evolutionary Prediction Games

### v1

@@ 5.4 Stabilizing Coexistence Equilibria
We have thus far considered retraining with the objective of maximizing accuracy at each time step on the current distribution. But if the learner is aware of population dynamics, then this can be leveraged in training. Our next result shows that it is possible to turn an unstable equilibrium into a stable one by simple reweighing of the accuracy objective:
Proposition 5.4.
Let $\mathcal{A}({\bm{p}})$ be an optimal algorithm with equilibrium ${\bm{p}}^{*}$. If ${\bm{p}}^{*}$ has full support (and thus is unstable), then it becomes stable under $\mathcal{A}^{\prime}({\bm{p}})=\mathcal{A}(2{\bm{p}}^{*}-{\bm{p}})$.
Proof in Section B.8. The idea is that $\mathcal{A}^{\prime}$ trains ‘as if’ the state is $2{\bm{p}}^{*}-{\bm{p}}$, which can be achieved by weighing examples from each group $k$ by $w_{k}=\frac{p_{k}}{2p^{*}_{k}-p_{k}}$. This inverts the natural tendency of dynamics to push away from the unstable equilibrium, and instead pull towards it.
6 Empirical Evaluation
Figure 5: Empirical evaluation results. (Left) Evolutionary prediction game induced by CIFAR-10 with Resnet-9 and groups representing horizontal flips. The game has beneficial unstable coexistence (Section 6.1). (Center) Simulated stabilization of replicator dynamics on the CIFAR-10 game (Section 5.4). (Right) Game induced by MNIST with label noise, showing stable beneficial coexistence (Section 6.2).
We now turn to explore evolutionary prediction games empirically using real data and simulated dynamics.
6.1 Beneficial Coexistence Through Data Augmentation
Our first empirical result explores an avenue for beneficial coexistence inspired by data augmentation techniques. Data augmentations enrich the training set with class-preserving transformations of inputs. These serve as a proxy for additional data that is representative of the task but is unbounded to the biases specific to the observed data distribution (e.g., camera angle, source of lighting). Such relations can occur naturally: for example, medical images from different hospitals may differ in imaging device calibration, and pooling them together can provide robustness.
@@ 6.1 Beneficial Coexistence Through Data Augmentation
We use the CIFAR-10 image recognition dataset (Krizhevsky, 2009), which consists of 60000 32x32 color images in 10 classes, with 6000 images per class. As horizontal image flips are considered class-preserving for this dataset, group $A$ consists of images sampled directly from CIFAR-10, and group $B$ consists of images under horizontal flip. We use a ResNet-9 network for prediction (He et al., 2016), and train it using the ffcv framework with default optimization parameters (Leclerc et al., 2023). For each ${\bm{p}}$ we measure the prediction model’s accuracy on the original images from the CIFAR test set (representing $\mathrm{acc}_{A}(h_{\bm{p}})$), and on their flipped counterparts (representing $\mathrm{acc}_{B}(h_{\bm{p}})$). For stabilization, we use the method described in Proposition 5.4 with ${\bm{p}}^{*}=(0.5,0.5)$. We simulate the evolutionary process using discrete replicator dynamics, and use linear interpolation due determine intermediate fitness values.
@@ Results.
A graphical representation of the estimated evolutionary prediction game is presented in Fig. 5 (Left). The game has three equilibria: Two stable equilibria with single-group dominance ($92.6\%{}$ accuracy), and an unstable coexistence equilibrium ($93.5\%{}$ accuracy). Fig. 5 (Center) shows how stabilization induces stable beneficial coexistence.
6.2 Stable Coexistence Through Overparameterization
Our second empirical result explores an avenue for stable coexistence driven by properties of overparametrized neural networks. Modern neural networks are frequently overparameterized, and are known to achieve strong test-time performance despite memorizing the training data (e.g., Zhang et al., 2021). Recent research attributes this phenomenon to the implicit biases introduced by the training process (e.g., Vardi, 2023). Building on the ideas presented in Theorem 5.3, we demonstrate that stable coexistence can emerge when an overparameterized neural network is trained on a multi-class classification task with label noise.
@@ 6.2 Stable Coexistence Through Overparameterization
We use the MNIST dataset for digit recognition (LeCun, 1998), which consists of 70000 28x28 grayscale images of handwritten digits across 10 classes. Each class is split unevenly between the two groups: Group $A$ is biased towards even digits ${\left\{{0,2,4,6,8}\right\}}$, and group $B$ is biased towards odd digits, both with a 4:1 imbalance. For each group, we introduce label noise to the majority classes, mapping the true label of each digit $d$ to the next digit with same parity with probability $0.2$ (i.e. for group $A$, label noising stochastically maps $0\mapsto 2$, $2\mapsto 4$, etc.). We train a convolutional neural network for 200 epochs using stochastic gradient decent with momentum. The training and test sets are split independently, and for each ${\bm{p}}$ we measure the prediction model’s accuracy on both splits of the test sets (representing $\mathrm{acc}_{A}(h_{\bm{p}})$ and $\mathrm{acc}_{B}(h_{\bm{p}})$). Under the given parameters, the label noising process leads to an accuracy upper bound of $84\%$ assuming perfect recognition.
@@ Results.
A graphical representation of the estimated evolutionary prediction game is presented in Figure 5 (Right). The game has a single beneficial coexistence equilibrium with accuracy $80.4\%{}$ (c.f. the $84\%$ test accuracy upper bound due to label noise). In addition, we observe an average training accuracy of $98.7\%{}$, indicating memorization.
@@ 7 Discussion
Impact statement
@@
Appendix A Further Discussion of Related Work
A.1 Relation to Performative Prediction
Performative prediction (Perdomo et al., 2020) is a general framework for modeling the effects of predictive models on the data distributions they are tested and trained on. In particular, stateful performative prediction (Brown et al., 2022) generalizes the original framework to model stateful performative dynamics. In the stateful framework, the performative effects of a predictor $h\in\mathcal{H}$ are formally modeled using a stateful transition function $D_{t+1}=T(D_{t},h)$, where $D_{t},D_{t+1}\in{\Delta\left({{\mathcal{X}\times\mathcal{Y}}}\right)}$ are the data distributions at times $t,t+1$, and $h\in\mathcal{H}$ is the predictor being deployed. We note that the space of feature-label distributions is infinite-dimensional under the common assumption $\mathcal{X}\subseteq\mathbb{R}^{d}$. Additionally, we note that the space of classifiers $\mathcal{H}$ is high-dimensional when the hypotheses class $\mathcal{H}$ is complex (e.g., a class of deep neural networks). Thus, the general framework allows for transition operators which are infinite-dimensional (e.g., as noted by Conger et al., 2024).
@@ A.1 Relation to Performative Prediction
In the context of performative prediction, our setting can be seen as a special case of stateful performative prediction, where the transition operator $T$ operates within a low-dimensional subspace of all possible feature label-distributions (mixtures of fixed distributions), and depends on predictors through low-dimensional statistics (their marginal accuracy on each group). Formally, given a collection of group distributions $D_{1},\dots,D_{K}\in{\Delta\left({{\mathcal{X}\times\mathcal{Y}}}\right)}$, denote the space of mixtures by $\Delta_{\mathrm{mix}}={\left\{{\sum_{k}\alpha_{k}D_{k}\mid\alpha\in{\Delta\left({[K]}\right)}}\right\}}$, and note that $\Delta_{\mathrm{mix}}\subseteq{\Delta\left({{\mathcal{X}\times\mathcal{Y}}}\right)}$. By definition, any mixture $D\in\Delta_{\mathrm{mix}}$ can be represented using a $(K-1)$-dimensional vector ${\bm{p}}\in{\Delta\left({[K]}\right)}$. Additionally, as performative response is assumed to depend on predictor accuracy, the transition operator depends on the classifier $h$ through a $K$-dimensional vector of accuracies $\mathbf{acc}(h)=\left(\mathrm{acc}_{1}(h),\dots,\mathrm{acc}_{K}(h)\right)\in\left[0,1\right]^{K}$. Therefore, transition operators induced by the evolutionary model can be seen as mappings ${\bm{p}}^{\prime}=T({\bm{p}},\mathbf{acc}(h))$ from a $(2K-1)$-dimensional space to a $(K-1)$-dimensional space.
Appendix B Deferred Proofs
B.1 Preliminaries

### v2

@@ 5.2 Stabilizing Coexistence Equilibria
Thus far we have considered evolutionary prediction games in which learning pursues the conventional objective of maximizing accuracy at each timestep on the current distribution. But if the learner is aware of population dynamics and of how the learned classifier may influence outcomes, then it makes sense to consider learning algorithms that take this into account. The question of course is how.
Here we focus on a particular aspect of this general question that arises from our last result in Sec. 4, namely: given a system with a desirable but unstable mixed equilibrium ${\bm{p}}^{*}$, how do we stabilize it? We propose a conceptually simple solution that works by reweighing examples in the standard accuracy objective in a way that accounts for dynamcis. The main idea is to invert the natural tendency of dynamics to push away from the unstable equilibrium, and instead pull towards it, by training ‘as if’ the distribution was at a different state ${\bm{p}}^{\prime}$ than its actual current state ${\bm{p}}$:
Proposition 1.
Let $\mathcal{A}^{\mathrm{opt}}({\bm{p}})$ be an oracle algorithm with equilibrium ${\bm{p}}^{*}$. If ${\bm{p}}^{*}$ has full support (and thus is unstable), then it becomes stable under $\mathcal{A}^{\prime}({\bm{p}})=\mathcal{A}^{\mathrm{opt}}(2{\bm{p}}^{*}-{\bm{p}})$.
Proof in Section D.9. The idea is that $\mathcal{A}^{\prime}$ trains ‘as if’ the state is $2{\bm{p}}^{*}-{\bm{p}}$, which can be achieved by weighing examples from each group $k$ by $w_{k}=\frac{p_{k}}{2p^{*}_{k}-p_{k}}$. This inverts the natural tendency of dynamics to push away from the unstable equilibrium, and instead pull towards it. The statement regards optimal classifiers and requires knowledge of ${\bm{p}}^{*}$, but in practice we can set $w$ according to an estimated $\tilde{{\bm{p}}}$. We further explore this approach in Section 6.2.
6 Experiments
Figure 4: Empirical evaluation results. (Left) Evolutionary prediction game induced by CIFAR-10 with Resnet-9 and groups representing horizontal flips. The game has unstable mutualism (Section 6.2). (Center) Simulated stabilization of replicator dynamics on the CIFAR-10 game (Section 5.2). (Right) Game induced by MNIST with label noise, showing stable mutualism (Section 6.1).
We now turn to explore evolutionary prediction games empirically using real data and simulated dynamics. Our goal here is to demonstrate how some of the principles we have discussed so far apply in more realistic settings. We include two experiments, designed to complement our results from sections 5.1 and 5.2, respectively. The first experiment shows how mutualism can arise organically due to group balancing, here from complementarities in label noise. The second demonstrates how stabilization can improve both overall and per-group accuracies by facilitating coexistence.
6.1 Stable Coexistence From Overparameterization
Modern neural networks are frequently overparameterized, and are known to achieve strong test-time performance despite memorizing the training data [e.g., 54]. Recent research attributes this phenomenon to the implicit biases introduced by the training process [e.g., 50]. Building on these ideas, here we demonstrate that stable coexistence can emerge when an overparameterized neural network is trained on data in which groups have complementing label noise. Our construction considers a multi-class settings where each group includes a majority of noisy examples from some class, and a minority of clean examples from other classes. This aims to captures settings in which user groups (e.g., by country) differ in the marginal distribution $p(y)$ over classes (e.g., spoken language), and where the dominant class is more likely to be misclassified (e.g., due to many accents).
@@ 6.1 Stable Coexistence From Overparameterization
We use the MNIST dataset for digit recognition [30], which consists of 70,000 28x28 grayscale images of handwritten digits across 10 classes. Each class is split unevenly between the two groups: Group $A$ is biased towards even digits ${\left\{{0,2,4,6,8}\right\}}$, and group $B$ is biased towards odd digits, both with a 4:1 imbalance. For each group, we introduce label noise to the majority classes, mapping the true label of each digit $d$ to the next digit with same parity with probability $0.2$ (i.e. for group $A$, label noising stochastically maps $0\mapsto 2$, $2\mapsto 4$, etc.). We train a convolutional neural network for 200 epochs using stochastic gradient decent with momentum. The training and test sets are split independently, and for each ${\bm{p}}$ we measure the prediction model’s accuracy on both splits of the test sets (representing $\mathrm{acc}_{A}(h_{\bm{p}})$ and $\mathrm{acc}_{B}(h_{\bm{p}})$). Under the given parameters, the label noising process leads to an accuracy upper bound of $84\%$ assuming perfect recognition.
@@ Results.
Figure 4 (Right) shows the estimated evolutionary prediction game, averaged across 505050{}50 repetitions. The game has a stable coexistence equilibrium, with $80.4\pm 0.2\%{}$ test accuracy (vs. the theoretical upper bound of $84\%$ due to label noise). This indicates that groups naturally balance each other in this setting. Note that training accuracy is $98.7\pm 0.1\%{}$, suggesting that interpolation is indeed the likely phenomenon at play.
6.2 Mutualistic Coexistence From Data Augmentation
Data augmentation is a method for enriching the training set with class-preserving transformations of inputs (e.g., camera angle, source of lighting) [e.g., 46]. Augmentations are typically considered as artificial constructs, but there are also cases where they emerge naturally (e.g., medical images from different hospitals or imaging devices). Here we show that when different sources correspond to different groups, data pooling can serve to foster beneficial coexistence in which all groups gain from the existence of others. This suggests that augmentations remain effective in the long run.
@@ 6.2 Mutualistic Coexistence From Data Augmentation
We use the CIFAR-10 image recognition dataset [27], which consists of 60,000 32x32 color images in 10 classes, with 6,000 images per class. As horizontal image flips are considered class-preserving for this dataset, group $A$ consists of images sampled directly from CIFAR-10, and group $B$ consists of images under horizontal flip. We use a ResNet-9 network for prediction [20], and train it using the ffcv framework with default optimization parameters [29]. For each ${\bm{p}}$, we measure the prediction model’s accuracy on the original images from the CIFAR test set (representing $\mathrm{acc}_{A}(h_{\bm{p}})$), and on their flipped counterparts (representing $\mathrm{acc}_{B}(h_{\bm{p}})$). For stabilization, we use the method described in Section 5.2 with ${\bm{p}}^{*}=(0.5,0.5)$. We simulate the evolutionary process using discrete replicator dynamics, and use linear interpolation to determine intermediate fitness values.
@@ Results.
Figure 4 (Left) presented the estimated evolutionary prediction game, averaged across 202020{}20 repetitions. The game has three equilibria: Two stable equilibria with single-group dominance ($92.6\pm 0.1\%{}$ accuracy), and an unstable coexistence equilibrium ($93.5\pm 0.1\%{}$ accuracy). The mixed equilibrium is optimal for both groups in terms of accuracy, but requires stabilization to become feasible. Figure 4 (Center) shows how our algorithm from Section 5.2 is able to attain stable beneficial coexistence, despite using an empirical proxy objective.
@@ 7 Discussion
This work introduces evolutionary prediction games, a unified framework for analyzing self-selection feedback loops between learning algorithms and user populations. Our analysis reveals a gap: in idealized settings, repeated training drives competitive exclusion, whereas practical constraints enable a rich variety of long-term outcomes, including stable coexistence and mutualistic interactions.
@@
Acknowledgements
The authors would like to thank Mor Nitzan, Oren Kolodny, Ariel Procaccia, Fernando P. Santos, Jill-Jênn Vie, Evyatar Sabag, and Anonymous Reviewers for their insightful remarks and valuable suggestions. Nir Rosenfeld is supported by the Israel Science Foundation grant no. 278/22. Eden Saig is supported by the Israel Council for Higher Education PBC scholarship for Ph.D. students in data science.
Appendix A Broader Impact and Limitations
@@
Appendix B Further Discussion of Related Work
B.1 Performative Prediction
Stateful performative prediction.
Performative prediction [39] is a framework for modeling the effects of predictive models on the data distributions they are tested and trained on (see [18] for a recent survey). Dynamics induced by evolutionary prediction games are a special case of the stateful performative prediction framework [3, 41], which models stateful feedback loops. The performative effects of a predictor $h\in\mathcal{H}$ are formally modeled using a transition map $D_{t+1}=\mathcal{T}(D_{t},h)$, where $D_{t},D_{t+1}\in{\Delta\left({{\mathcal{X}\times\mathcal{Y}}}\right)}$ are the data distributions at times $t,t+1$, and $h\in\mathcal{H}$ is the predictor being deployed. The fixed points of the dynamics induced by evolutionary prediction games (i.e. equilibrium states ${\bm{p}}^{*}$), correspond to fixed point distributions as defined in [3, Definition 4], as they satisfy $D_{{\bm{p}}^{*}}=\mathcal{T}(D_{{\bm{p}}^{*}},h_{{\bm{p}}^{*}})$, where $h_{{\bm{p}}^{*}}$ is the classifier trained on $D_{{\bm{p}}^{*}}$. Moreover, we note that the notion of evolutionary stability is stronger than the notion of fixed point distributions, as evolutionarily stable points also remain stable under small perturbations around ${\bm{p}}^{*}$.
Structural assumptions reduce effective dimensionality.
Compared to the general performative prediction framework, natural selection dynamics induce transition maps which are lower-dimensional, and thus more amenable for analysis. The general space of feature-label distributions is infinite-dimensional under the common assumption $\mathcal{X}\subseteq\mathbb{R}^{d}$, and the space of classifiers $\mathcal{H}$ is high-dimensional when the hypotheses class $\mathcal{H}$ is complex (e.g., deep neural networks). Thus, the general performative prediction framework allows for transition operators which are infinite-dimensional and notoriously challenging to analyze (e.g., as noted by Conger et al. [8]). In contrast, in dynamics induced by evolutionary prediction games, the stateful transition operator $\mathcal{T}$ operates within a low-dimensional subspace of all possible feature label-distributions (mixtures of fixed distributions), and depends on predictors through low-dimensional statistics (their marginal accuracy on each group).
@@ B.1 Performative Prediction
Formally, given a collection of group distributions $D_{1},\dots,D_{K}\in{\Delta\left({{\mathcal{X}\times\mathcal{Y}}}\right)}$, denote the space of mixtures by $\Delta_{\mathrm{mix}}={\left\{{\sum_{k}p_{k}D_{k}\mid{\bm{p}}\in\Delta^{K}}\right\}}$, and note that $\Delta_{\mathrm{mix}}\subseteq{\Delta\left({{\mathcal{X}\times\mathcal{Y}}}\right)}$. Any mixture $D\in\Delta_{\mathrm{mix}}$ can be represented by a point on the $(K-1)$-dimensional simplex ${\bm{p}}\in\Delta^{K}$. Additionally, as performative response depends on predictor accuracy, the transition operator depends on the classifier $h$ through a $K$-dimensional vector of accuracies $\mathbf{acc}(h)=\left(\mathrm{acc}_{1}(h),\dots,\mathrm{acc}_{K}(h)\right)\in\left[0,1\right]^{K}$. Therefore, transition operators induced by the evolutionary model can be viewed as lower-dimensional mappings ${\bm{p}}^{\prime}=\mathcal{T}({\bm{p}},\mathbf{acc}(h))$ from a $(2K-1)$-dimensional space to a $(K-1)$-dimensional space.
Further relations to performative risk and forecasting.
The gaps between the average accuracies of equilibrium points relates to the distinction between performative optimality and performative stability [e.g., 35]. Additionally, convex functions over the simplex (and their sub-gradients) emerge as a key component in our analysis of oracle classifiers (Section 4), and also appears as a key component in the analysis of proper scoring rules and their performative counterparts [36, 6] through the fundamental characterization theorem [15, 44]. From this perspective, interpreting the convex potential of oracle-learner games as a generator of proper scoring rules is an intriguing direction for future work.
B.2 Long-term Fairness
The long-term fairness literature studies the long-term implications of population dynamics on algorithmic fairness (see [16] for a recent survey). In particular, several works in this area also consider different variations of replicator (and replicator-like) dynamics as a model for user response [e.g., 19, 40, 53, 11]. Perhaps closest to our approach is Hashimoto et al. [19], who study worst-group accuracy dynamics under repeated training in a mixture population that evolves via stochastic replicator-like dynamics supplemented by a strong, balanced influx of new users that join the system regardless of its performance (parameterized by $b_{k}$ in their model). While their stability guarantees rely on the assumption of this strong user flow, our results offer a complementary view, showing that stable coexistence may emerge even without an exogenous user flow, under standard empirical risk minimization, and without compromising worst-group accuracy (see Section 5). We also extend their ERM instability results to a wider family of behavioral dynamics, and establish global convergence guarantees, rather than local (see Section 4, and Appendix C).
B.3 Dynamics-Aware Learners
Finally, our framework models the learning algorithm as an entity which is only informed by the current state of population $h\sim\mathcal{A}({\bm{p}})$. While still common in many practical scenarios (see, e.g., [12, 24, 25]), we also note that various practical systems already optimize for long-term objectives. Examples includes learners that reduce hidden incentives to modify the population of users [28, Sec. 7.3], adaption to shifting user preferences [23, 4], or directly maximizing long-term engagement via reinforcement learning (e.g., [1]). While our focus here is on local learning rules, analyzing dynamics-aware learners from an evolutionary perspective is an intriguing direction for future inquiry.
Appendix C Microfoundations
Natural selection dynamics appear in a variety of settings. This appendix offers a representative (though not exhaustive) set of examples showing how local, myopic user behaviors aggregate into population-level dynamics that satisfy our core assumptions. Using online social platforms as a running example, we show how different behavioral patterns (e.g., invitation, imitation, resource allocation) and different types group affiliations (e.g., social, behavioral, demographic) may give rise to natural-selection dynamics in various parts of the ecosystem. We note that similar microfoundations can be identified across a wide range of domains, such as finance, healthcare, and education.
C.1 Reproduction
Consider a recommendation system based on prediction. Accurate recommendations motivate users to invite peers from their social group, while the presence of alternatives creates a risk of dropout.
Formally, assume that users from $K$ distinct groups interact with a prediction system over time, and denote by $N_{k}$ the number of users in group $k\in[K]$. On average, users interact with the system $\lambda>0$ times per time step $t=1,2,\dots$ (e.g., three times per day), and at each interaction they receive a prediction (e.g., content relevance). If a prediction is accurate, users invite a peer from their group with probability $\alpha$. Independently, a user may leave the system with probability $\beta$ (e.g., due to friction or better alternatives). In expectation, the size of each group is multiplied by the following factor:
$N_{k}^{t+1}$ $=\left(1+\lambda\left(\alpha\cdot\mathrm{acc}_{k}(h^{t})-\beta\right)\right)N_{k}^{t}$
In the continuous-time limit, users interact with the system $\lambda\mathrm{d}t$ times per step, and $\mathrm{d}t\to 0$. The expected growth approaches the differential equation:
$\dot{N}_{k}=\lambda(\alpha\cdot\mathrm{acc}_{k}(h)-\beta)N_{k}$
Where $\dot{N}_{k}=\tfrac{\mathrm{d}}{\mathrm{d}t}N_{k}$ is the time derivative of $N_{k}$. Denote $F_{k}=\lambda(\alpha\cdot\mathrm{acc}_{k}(h)-\beta)$, such that $\dot{N}_{k}=F_{k}N_{k}$. Note that the growth multiplier $F_{k}$ is time-dependent when the classifier $h$ is retrained.
For group-level dynamics, denote by $N=\sum_{k}N_{k}$ the total number of users, and by $p_{k}=\tfrac{N_{k}}{N}$ the proportion of group $k$. The time derivative of $p_{k}$ is:
$\dot{p}_{k}$ $=\frac{\mathrm{d}}{\mathrm{d}t}\frac{N_{k}}{N}$
$=\frac{\dot{N}_{k}N-\dot{N}N_{k}}{N^{2}}$
$=\underbrace{\frac{N_{k}}{N}}_{=p_{k}}\bigg{(}\underbrace{\frac{\dot{N}_{k}}{N_{k}}}_{\dot{N}_{k}=F_{k}N_{k}}-\underbrace{\frac{\dot{N}}{N}}_{\dot{N}=\sum F_{k}N_{k}}\bigg{)}$
$=p_{k}\left(F_{k}-\sum_{k^{\prime}\in[K]}p_{k^{\prime}}F_{k^{\prime}}\right)$
Thus, the aggregate dynamics of group proportions ${\bm{p}}$ satisfy the replicator equation [48], which satisfies our core assumptions. The fitness function in this case is $F_{k}({\bm{p}})=\lambda\alpha\cdot\mathrm{acc}_{k}(h_{\bm{p}})-\lambda\beta$.
C.2 Imitation
Consider a media platform fostering content creators, whose profit depends on the extent to which their content reaches their target audience through accurate predictions. A successful approach (e.g., content style, advertising strategy) of one creator is likely to be mimicked by others.
Formally, suppose there a large population of content producers aiming to market some good. Each producer has a content production strategy $k$ (e.g., short versus long videos), the price of the good is $r$, and the cost of production under strategy $k$ is $c_{k}$ (e.g., longer videos are typically more costly to produce). A prediction algorithm observes content and makes recommendation decisions to content consumers. A prediction is considered accurate if the content is relevant to the consumer, and each successful prediction yields a conversion (e.g., sale) with probability $\alpha$. Denote by $x\sim D_{k}$ a content item produced by strategy $k$, and its corresponding label by $y$. The expected profit of a producer using strategy $k$ is:
$F_{k}=r\cdot\alpha\mathbb{P}_{D_{k}}{\left[{h(x)=y}\right]}-c_{k}$
At each step, a content creator observes another creator’s content (e.g., on social media). If their own strategy is less successful than the one they observe, they adopt the other’s strategy with probability proportional to the difference in utilities. Denote the total size of the population by $N$ (here assumed to be large but fixed), the number of creators who adopt strategy $k$ by $N_{k}$, and the proportion of group $k$ by $p_{k}=N_{k}/N$. Sandholm [43, Chapter 4] shows that this form of pairwise proportional imitation converges in the large-population limit to the replicator dynamics [48], which satisfy our core assumptions.
C.3 Online resource allocation
Natural-selection dynamics between groups of users can also arise from interaction between prediction algorithms in multi-stage pipelines. Consider a firm that acquires users though $K$ marketing campaigns, each targeting a different demographic. The firm uses a classifier to personalize the experience of incoming users, and each accurate prediction yields a conversion with probability $\alpha$. Campaign $k$ therefore generates expected utility:
$F_{k}=r\cdot\alpha\cdot\mathrm{acc}_{k}(h)-c_{k}$
where $r$ is the revenue per conversion, and $c_{k}$ is the average cost of user acquisition for campaign $k$. The firm has a fixed advertising budget $B$, which it dynamically allocates across marketing campaigns. Denote the spend on campaign $k$ by $p_{k}B$, and assuming the firm reassigns spend according to a multiplicative weights update rule [13] for dynamic allocation, we obtain:
$p_{k}^{t+1}=p_{k}\frac{\beta^{F_{k}}}{\sum_{k^{\prime}}\beta^{F_{k^{\prime}}}}$
where $\beta>1$ is a hyper-parameter of the MW algorithm. As $\beta\to 1$, dynamics converge towards the replicator dynamics (cf. Equation 17), which satisfies our core assumptions.
Appendix D Deferred Proofs
D.1 Preliminaries
Figure 5: Graphical representation of two-group evolutionary prediction games (Definition 1). Groups are denoted by ${\left\{{A,B}\right\}}$. Solid lines describe expected group accuracies for each population state ${\bm{p}}$, dotted line describes population accuracy $\mathrm{acc}_{\bm{p}}(h)$, dots indicate Nash equilibria, and x-axis arrows indicate evolutionary stability. Schematic plots illustrate a selection of possible scenarios: (Left) Competitive exclusion. There are two stable equilibria, each dominated by a single group, with locally maximal population accuracy. Additionally, there is an unstable coexistence equilibrium with lower overall accuracy. (Center Left) Stable mutualistic coexistence. There is a stable coexistence equilibrium, where the accuracy of both groups is maximized. (Center Right) Unstable mutualisitic coexistence. Coexistence has maximal population accuracy, but it is unstable. (Right) No coexistence equilibria. There is a single equilibrium dominated by a single group.